# Symmetric Mass Generation in a Bilayer Honeycomb Lattice with $\mathrm{SU}(2)\times\mathrm{SU}(2)\times\mathrm{SU}(2)/\mathbb{Z}_{2}$ Symmetry

### 18 Mar 2026 v1

@@ Abstract
Symmetric mass generation (SMG) is a mechanism for generating mass gaps in fermionic systems without breaking any symmetries or developing topological order, challenging the conventional Landau paradigm. In this Letter, we provide numerically exact evidence for SMG in (2+1) dimensions through large-scale determinant quantum Monte Carlo (DQMC) simulations of a bilayer honeycomb lattice model with $\mathrm{SU}(2)\times\mathrm{SU}(2)\times\mathrm{SU}(2)/\mathbb{Z}_{2}$ symmetry. We observe the simultaneous opening of single-particle and bosonic gaps at a critical coupling $J_{c}\approx 2.6$ with correlation length exponent $\nu=1.14(2)$, while an exhaustive search over all 19 symmetry-inequivalent fermion bilinear order parameters confirms the absence of any symmetry breaking. We estimate the fermion anomalous dimension to be $\eta_{\psi}=0.071(1)$, which deviates significantly from the large-$N$ prediction ($\eta_{\psi}\approx 0.595$) and variational Monte Carlo estimates ($\eta_{\psi}\approx 0.62$), pointing to a distinct universality class. By contrasting with a related $\mathrm{Spin}(5)\times\mathrm{U}(1)/\mathbb{Z}_{2}$ model that develops an intermediate excitonic phase, we demonstrate the crucial role of pure non-Abelian symmetry in prohibiting bilinear condensates and enforcing the direct SMG transition.
@@
Recent theoretical developments have revealed the possibility of generating mass gaps without breaking any symmetries or developing topological order, a phenomenon known as symmetric mass generation (SMG) [5, 6, 7]. This mechanism represents a significant departure from the Landau paradigm, which classifies phases based on their symmetry-breaking patterns. In SMG, the gap is not a single-particle property but emerges from many-body correlations: the mass gap equals the inverse of the many-body correlation length and vanishes only at the quantum critical point. SMG thus allows for direct, continuous transitions from gapless to gapped phases while preserving all symmetries, suggesting exotic quantum critical points beyond conventional field-theoretic descriptions.
@@
Various models exhibiting SMG have been proposed across different dimensions. In (0+1)D, the Majorana fermion model studied by Fidkowski and Kitaev [9, 10] demonstrated how interactions can modify the classification of topological phases. In (1+1)D, the 3-4-5-0 chiral fermion model [11] provided a concrete lattice framework for regularizing chiral matter. For (2+1)D, several models have been investigated, including bilayer honeycomb lattice models [12, 13, 14, 15] and staggered fermion models [16, 17, 18]. In (3+1)D, chiral fermion models have been explored for their potential to exhibit SMG [7, 19].
Despite these theoretical advances, unbiased numerical evidence for SMG in (2+1)D has remained elusive. Especially, the unbiased estimation of the fermion’s scaling dimension remains to be tested. High symmetry is a key ingredient, as it introduces large quantum fluctuations that can prevent symmetry breaking, but it alone does not guarantee a direct transition to a gapped phase. Indeed, recent variational Monte Carlo (VMC) studies of bilayer honeycomb models [14] could not fully settle the question due to the uncontrolled approximations inherent in the variational approach. Rigorous, sign-problem-free quantum Monte Carlo calculations are essential to determine whether a given model truly exhibits SMG or instead develops intermediate symmetry-breaking phases.
In this Letter, we employ large-scale determinant quantum Monte Carlo (DQMC) simulations on the bilayer honeycomb lattice model proposed by Hou and You [14] [Fig. 1(a)], with $\mathrm{SU}(2)\times\mathrm{SU}(2)\times\mathrm{SU}(2)/\mathbb{Z}_{2}$ (abbreviated as $\mathrm{SU}(2)^{3}/\mathbb{Z}_{2}$ in the following) symmetry. Our results provide the first numerically exact confirmation of a direct, continuous SMG transition in (2+1) dimensions and yield a fermion anomalous dimension $\eta_{\psi}=0.071(1)$ that differs significantly from both VMC and large-$N$ predictions, pointing to a distinct universality class. By contrasting with a related $\mathrm{Spin}(5)\times\mathrm{U}(1)/\mathbb{Z}_{2}$ model that develops an intermediate excitonic phase (see the Supplemental Material (SM) [20]), we demonstrate the essential role of pure non-Abelian symmetry in enforcing the direct SMG transition. The systematic framework for identifying symmetries and classifying all fermion bilinear order parameters is developed in a companion paper [21].
@@
A hallmark of the SMG mechanism is the absence of spontaneous symmetry breaking (SSB) in the gapped phase. To verify this, we performed an exhaustive group-theoretical classification of all possible fermion bilinear order parameters locally within the unit cell. Within the systematic framework developed in the companion paper [21], the 119 nontrivial Majorana bilinears decompose into 19 irreducible multiplets under the full symmetry group, which is a $\mathbb{Z}_{2}^{s}\times\mathbb{Z}_{2}^{l}$ extension of $\mathrm{SU}(2)^{3}/\mathbb{Z}_{2}$, where $\mathbb{Z}_{2}^{s}$ and $\mathbb{Z}_{2}^{l}$ denote sublattice exchange and layer exchange, respectively. (see the SM [20] for the complete list). We computed the structure factors for each multiplet across the transition. Remarkably, finite-size scaling analysis confirms that all 19 order parameters extrapolate to zero in the thermodynamic limit for all values of $J$ (for more details, see the SM [20]). Fig. 3 shows that conventional ferromagnetic (FM), charge density wave (CDW), spin density wave (SDW), and EC/SC order are explicitly ruled out as representative examples. The absence of any SSB supports the realization of a featureless, symmetric insulator.
@@
The critical exponents extracted from our study challenge current effective field theory descriptions. While the transition shares qualitative features with the proposed fDQCP scenario, the quantitative mismatch—our $\eta_{\psi}=0.071(1)$ versus the VMC value $\sim 0.62$ [14] and the large-$N$ prediction $\sim 0.595$ [24]—points toward a more complex low-energy effective theory. It is instructive to contrast our findings with a closely related bilayer honeycomb model possessing $\mathrm{Spin}(5)\times\mathrm{U}(1)/\mathbb{Z}_{2}$ symmetry (see the SM [20] for a detailed analysis). In that system, we found no direct SMG transition; instead, an intermediate phase characterized by inter-layer excitonic order with $\mathrm{U}(1)$ symmetry breaking emerges between the semimetal and the trivial insulator (the first transition belongs to the Gross-Neveu-XY universality class and the second to the (2+1)D XY class). This comparison highlights that the pure non-Abelian symmetry of the $\mathrm{SU}(2)^{3}/\mathbb{Z}_{2}$ model is essential for prohibiting bilinear condensates and enforcing the direct SMG transition: the absence of any $\mathrm{U}(1)$ factor prevents the formation of excitonic order that would otherwise intervene.

### 17 Apr 2026 v2

@@ Abstract
A central question beyond the Landau paradigm is the non-perturbative critical theory of the symmetric mass generation (SMG) transition, where strong interactions gap Dirac fermions in (2+1) dimensions without triggering spontaneous symmetry breaking or topological order. While previous studies have already provided evidence for direct SMG transitions in (2+1) dimensions, the fermion scaling dimension—the key observable for distinguishing candidate critical theories—has not been determined in a controlled unbiased way. In this Letter, using large-scale determinant quantum Monte Carlo (DQMC) simulations of a bilayer honeycomb lattice model with $\mathrm{SU}(2)\times\mathrm{SU}(2)\times\mathrm{SU}(2)/\mathbb{Z}_{2}$ symmetry, we establish a direct continuous transition by observing the simultaneous opening of single-particle and bosonic gaps at a critical coupling $J_{c}\approx 2.6$ with correlation length exponent $\nu=1.14(2)$, while an exhaustive search over all 19 symmetry-inequivalent fermion bilinear order parameters confirms the absence of any symmetry breaking. We further obtain the first controlled unbiased estimate of the fermion anomalous dimension, $\eta_{\psi}=0.071(1)$, which deviates significantly from the large-$N$ prediction ($\eta_{\psi}\approx 0.595$) and variational Monte Carlo estimates ($\eta_{\psi}\approx 0.62$), thereby placing direct quantitative constraints on SMG criticality. By contrasting with a related $\mathrm{Spin}(5)\times\mathrm{U}(1)/\mathbb{Z}_{2}$ model that develops an intermediate excitonic phase, we show that pure non-Abelian symmetry plays a decisive role in stabilizing the direct SMG transition.
@@
Recent theoretical developments have revealed the possibility of generating mass gaps without breaking any symmetries or developing topological order, a phenomenon known as symmetric mass generation (SMG) [5, 6, 7, 8, 9, 10, 11]. The idea of gapping fermions through strong interactions without symmetry breaking has roots in early lattice studies of Yukawa models and mirror-fermion decoupling [12, 13, 14, 15, 16, 17, 18, 19, 20], where symmetric phases with massive fermions were identified at strong bare coupling, though the connection to anomaly cancellation and universal low-energy physics was clarified only later. This mechanism represents a significant departure from the Landau paradigm, which classifies phases based on their symmetry-breaking patterns.
@@
Various models exhibiting SMG have been proposed across different dimensions. In (0+1)D, the Majorana fermion model studied by Fidkowski and Kitaev [22, 23] demonstrated how interactions can modify the classification of topological phases. In (1+1)D, the 3-4-5-0 chiral fermion model [24] provided a concrete lattice framework for regularizing chiral matter. For (2+1)D, several models have been investigated, including bilayer honeycomb lattice models [25, 26, 27, 28] and staggered fermion models [29, 30, 31, 32]. In (3+1)D, chiral fermion models have been explored for their potential to exhibit SMG [7, 33].
Despite these theoretical advances, unbiased numerical evidence for SMG in (2+1)D has remained elusive. In particular, an unbiased determination of the fermion scaling dimension has remained out of reach. High symmetry is a key ingredient, as it introduces large quantum fluctuations that can prevent symmetry breaking, but it alone does not guarantee a direct transition to a gapped phase. Indeed, recent variational Monte Carlo (VMC) studies of bilayer honeycomb models [27] could not fully settle the question due to the uncontrolled approximations inherent in the variational approach. Rigorous, sign-problem-free quantum Monte Carlo calculations are essential to determine whether a given model truly exhibits SMG or instead develops intermediate symmetry-breaking phases.
In this Letter, we employ large-scale determinant quantum Monte Carlo (DQMC) simulations [34, 35, 36, 37, 38, 39, 40] on the bilayer honeycomb lattice model proposed by Hou and You [27] [Fig. 1(a)], with $\mathrm{SU}(2)\times\mathrm{SU}(2)\times\mathrm{SU}(2)/\mathbb{Z}_{2}$ (abbreviated as $\mathrm{SU}(2)^{3}/\mathbb{Z}_{2}$ in the following) symmetry. Our results provide the first numerically exact confirmation of a direct, continuous SMG transition in (2+1) dimensions and yield a fermion anomalous dimension $\eta_{\psi}=0.071(1)$ that differs significantly from both VMC and large-$N$ predictions, pointing to a distinct universality class. By contrasting with a related $\mathrm{Spin}(5)\times\mathrm{U}(1)/\mathbb{Z}_{2}$ model that develops an intermediate excitonic phase (see the Supplemental Material (SM) [41]), we demonstrate the essential role of pure non-Abelian symmetry in enforcing the direct SMG transition. The symmetry analysis and classification of all fermion bilinear order parameters used here are developed systematically in a companion paper [42].
@@
A hallmark of the SMG mechanism is the absence of spontaneous symmetry breaking (SSB) in the gapped phase. To verify this, we performed an exhaustive group-theoretical classification of all possible fermion bilinear order parameters locally within the unit cell. Using the framework developed in the companion paper [42], the 119 nontrivial Majorana bilinears decompose into 19 irreducible multiplets under the full symmetry group, which is a $\mathbb{Z}_{2}^{s}\times\mathbb{Z}_{2}^{l}$ extension of $\mathrm{SU}(2)^{3}/\mathbb{Z}_{2}$, where $\mathbb{Z}_{2}^{s}$ and $\mathbb{Z}_{2}^{l}$ denote sublattice exchange and layer exchange, respectively (see the SM [41] for the complete list). We computed the structure factors for each multiplet across the transition. Remarkably, finite-size scaling analysis confirms that all 19 order parameters extrapolate to zero in the thermodynamic limit for all values of $J$ (for more details, see the SM [41]). Fig. 3 shows that conventional ferromagnetic (FM), charge density wave (CDW), spin density wave (SDW), and EC/SC order are explicitly ruled out as representative examples. The absence of any SSB supports the realization of a featureless, symmetric insulator.
@@
The critical exponents extracted from our study challenge current effective field theory descriptions. While the transition shares qualitative features with the proposed fDQCP scenario, the quantitative mismatch—our $\eta_{\psi}=0.071(1)$ versus the VMC value $\sim 0.62$ [27] and the large-$N$ prediction $\sim 0.595$ [45]—points toward a more complex low-energy effective theory. It is instructive to contrast our findings with a closely related bilayer honeycomb model possessing $\mathrm{Spin}(5)\times\mathrm{U}(1)/\mathbb{Z}_{2}$ symmetry (see the SM [41] for a detailed analysis). In that system, we find no direct SMG transition; instead, an intermediate phase characterized by inter-layer excitonic order with $\mathrm{U}(1)$ symmetry breaking emerges between the semimetal and the trivial insulator (the first transition belongs to the Gross-Neveu-XY universality class and the second to the (2+1)D XY class). This comparison highlights that the pure non-Abelian symmetry of the $\mathrm{SU}(2)^{3}/\mathbb{Z}_{2}$ model is essential for prohibiting bilinear condensates and enforcing the direct SMG transition: the absence of any $\mathrm{U}(1)$ factor prevents the formation of excitonic order that would otherwise intervene.
@@ Acknowledgements.
Note added. We are aware a work on related topics that have appeared on arXiv after our posting [46]; their results have finite overlap with ours and seem consistent.